Minute of KAGRA commissioning meeting
- Date: 2018/7/3 11AM~12PM
- Attendance: Miyakawa, Aso, Miyoki, Enomonoto, Yokozawa
- eZuce: Suzuki, Shoda, Hirose, Izumi, Yamamoto(Toyama), Ohashi, Kimura, Tomaru, Inoue
40W laser and RIN measurement -> Tahara, Nakano
- Noise floor increase with power increment.
- Movement to Kamioka will be the end of July.
VCO noise measurement -> Sugimoto, Enomoto
- Any good idea to control only high frequency? (Aso)
- 40MHz less noise VCO can be used for also high frequency.
Green Fiber Noise Measurement -> Sugimoto, Enomoto
- Transmissivity was 60% first, then reduced to 40%.
- 5m fiber noise was measured and fitted, and will be scaled to 50m longer.
Geophone Noise Hunting -> Tanioka
- Circuit inside of pre-ampifier seems to be something wrong.
- LED light can be a noise. (Miyoki)
- Long cable could be noise, but after changing did not fix.
- Current monitor of HPCD can oscillate easily. At the coil output, no noise increment, no strange transfer function can be seen, but we did not check offset jump. We should do.
WAB alignment and the final saving-mode -> Akutsu, Simon
- Cable cleaning on BS rack will be done this weekend.
- pile of AA or HPCD can be hot.
Necessary number of High power coil driver -> Aso -> Current number of HPCD including 16 spares seems to be enough.
NAB PD cabling -> Aso
- Beam height brain storming
- ETMX and ITMX should be located at half low temperature stage and room temperature stage. It means that ~5mm off from the nominal position. Type A can compensate ~5mm.
Where is an actual beam hight? -> 30urad * ~10m -> not a big amount.
- Effect of window of gate valve?
- Next meeting: 7/17.